National Corporation needs to set a target price for its newly designed product M14–M16. The following data relate to this new product.

Per Unit Total
Direct materials $26
Direct labor $44
Variable manufacturing overhead $15
Fixed manufacturing overhead $1,377,000
Variable selling and administrative expenses $ 5
Fixed selling and administrative expenses $ 1,053,000


These costs are based on a budgeted volume of 81,000 units produced and sold each year. National uses cost-plus pricing methods to set its target selling price. The markup percentage on total unit cost is 40%.

Compute the total variable cost per unit, total fixed cost per unit, and total cost per unit for M14–M16.

Variable cost per unit $enter a dollar amount

Entry field with incorrect answer now contains modified data

Fixed cost per unit enter a dollar amount

Entry field with incorrect answer now contains modified data

Total cost per unit $enter a total of the two previous amounts

Entry field with incorrect answer now contains modified data

Compute the desired ROI per unit for M14–M16.

Desired ROI $enter the desired ROI per unit in dollars

Entry field with incorrect answer now contains modified data

per unit

Compute the target selling price for M14–M16.

Target selling price per unit $enter the target selling price per unit in dollars

Entry field with incorrect answer

Compute variable cost per unit, fixed cost per unit, and total cost per unit assuming that 60,750 M14–M16s are produced and sold during the year.

Variable cost per unit $enter a dollar amount

Entry field with incorrect answer

Fixed cost per unit enter a dollar amount

Entry field with incorrect answer

Total cost per unit $enter a total of the two previous amounts

Entry field with incorrect answer now contains modified data
Answer #1
Direct materials 26
Direct labor 44
Variable manufacturing overhead 15
Variable selling and administrative expenses 5
Variable cost per unit 90
Fixed cost per unit 30 =(1377000+1053000)/81000
Total cost per unit 120
Desired ROI 48 =120*40%
Total cost per unit 120
Add: Desired ROI per unit 48
Target selling price per unit 168
Variable cost per unit 90
Fixed cost per unit 40 =(1377000+1053000)/60750
Total cost per unit 130